# Toward General-Purpose Robots via Foundation Models: A Survey and Meta-Analysis

### v2

@@ 4.1.3 LLMs and VLMs in Action Generation
(a) Reward generation with LLM :
L2R [226]
(b) Locomotion control via action pattern
matching with LLM: SayTap [26]
(c) VLM for affordance function in motion
planning: VoxPoser [227]
Figure 8: LLM used in motion planning and action generation.
@@
4.1.4 Grounding in Robotics
Alongside the above dimensions, an equally crucial aspect is the concept of “grounding”. Grounding refers to the ability to associate contextual meaning with signals or symbols, such as the ability to tie a word to its manifestation in the world. Humans understand semantic concepts through both audio (words, tone) and visual signals (gestures, behaviors, body language). In the scope of this survey, grounding alludes to the process of aligning the abstract knowledge possessed by foundation models with the tangible, real-world specifics of robotics—ensuring that language-driven decisions correspond meaningfully with physical actions and environmental contexts. For instance, if we ask an LLM to generate a plan to find a pen in a specific house, without any environmental information, then this task is akin to a blind person reasoning about how to navigate in an unknown space, rendering the task nearly impossible to complete. Similarly, while an LLM may easily generate a plan to lift a chair by suggesting to ’grab the left handle with the left hand and the right handle with the right hand, then lift’, this plan becomes impractical if the real-world embodiment of the model is a typical robot equipped with only one arm. Since grounding is a large field in itself, we do not attempt to cover all of it; instead, we will address four concepts as shown in the Figure 9: (1) Grounding language to environments; (2) Grounding latent concepts to environments; (3) Grounding language to embodiments; and (4) Grounding latent concepts to embodiments.
Grounding Language to Environments
As discussed in the previous sections, specifically Sections 4.1.2 and 4.1.3, for the successful integration of LLMs or VLMs directly into robots, it is crucial to establish a connection between the language output and either skills or low-level actions. SayCan [24] learned a value function to score the joint likelihood between task-level plans and skills acquired through reinforcement learning or behavior cloning. ZSP [219] utilized semantic similarities between language instructions and skill names to ground robot skills, while ProgPrompt [221] and Code as Policy ground robot skills via program code synthesis. However, these skills are trained specifically for the environments in which they were tested. Therefore, we categorize all methods that associate language with skills as grounding to environments. Some research efforts, such as Grounded Decoding [233], attempt to address this by grounding foundation models to skills trained in a single environment, using small language-conditioned models trained in various environments as a probabilistic filter. Nonetheless, this approach is still limited as it presupposes the availability of grounded models suited for the environment in which it is embodied. In addition to skill grounding, work like Voxposer [227] have attempted to use LLMs to generate code for constructing value maps of the environment, which can then be planned on using existing planners. This represents a more general approach to integrating contextual information about the environment, compared to grounding to skill libraries.
Grounding Concepts to Environments
The term ‘concepts’ refers to the unified latent representations, derived from training with diverse input data. Approaches such as CLIP-Fields  [202], VLMap [201], and NLMaps [25] have endeavored to project CLIP visual representations and semantic label representations directly onto 3D point clouds. Beyond constructing explicit 3D maps, GLAM [234] proposes using reinforcement learning to ground language models with an internal map through interaction with environments. This approach demonstrated that LLMs can effectively function as Reinforcement Learning (RL) agents in textual environments. However, the challenge of generalizing these approaches to different environments and tasks, either by grounding concepts to point clouds or through implicit mapping, remains unresolved in current literature.
Grounding Language to Embodiments
We categorize language grounding to embodiments as necessitating a specific condition: it must be agnostic to different environments. This is akin to having a universal interface that translates language into actions. Initiatives like Prompt2Walk [235] and Saytap [26] have experimented with directly using LLMs to generate joint angles or foot patterns in the language space for locomotion control.
Grounding Concepts to Embodiments
@@ 4.1.4 Grounding in Robotics
Grounding concepts to embodiments involves directly anchoring foundation models to output joint torques, circumventing intermediary interfaces such as text. A notable example is Gato [30], which demonstrates versatility in tasks like playing Atari, captioning images, conversing, and manipulating objects with a robotic arm. Gato dynamically decides its output format—be it text, joint torques, button presses, or other tokens—based on the context of the task at hand. Another related development is RT-2 [28], which, despite specifying the end-effector space in textual form, is capable of directly generating executable commands for robot manipulator operation.
Figure 9: Grounding in robotics. We divide grounding techniques as grounding to environment and grounding to robotic embodiments.
In conclusion, the primary advantage of grounding language to environments and embodiments is the ease of implementing zero-shot learning without the need for additional training. However, significant drawbacks exist. For concepts not describable by language, such as the nuances of finger movements, grounding to embodiments may fail. Moreover, reliance on a fixed set of skill libraries limits adaptability to diverse environments. Consequently, direct grounding from the latent concept space appears to be a more viable solution. Approaches that utilize interaction data [234] or expert data [28] have both shown promising results in addressing these challenges.
@@ 4.2 Robotics Foundation Models (RFMs)
(a) Imitation learning: RT-1 [27]
(b) Reinforcement learning: Q-Transformer [195]
(c) Vision language pretraining: MOO [249]
Figure 10: Examples of Robotics Foundation Models (RFMs)
Robotic Foundation Models for: manipulation Robotics Foundation Models for: motion planning for navigation and multipurpose task
(d) Motion planning model: VINT [29]
(e) General purpose model: Gato [30]
(f) General purpose model: PaLM-E [31]
With the increasing amount of robotics datasets, containing state-action pairs from real robots, the class of Robotics Foundation Models (RFMs) have likewise become increasingly viable [28, 176, 29]. These models are characterized by the use of robotics data to train them, in order to solve robotics tasks. In this subsection, we summarize and discuss different types of RFMs. We will first introduce RFMs that can perform one set of tasks, within one of the robotic modules from Section 2.1, which is defined as single-purpose Robotic Foundation Models. For example, an RFM that can generate low-level actions to control the robots, or a model that can generate higher-level motion plans. We later introduce RFMs that can carry out tasks in multiple robotic modules, hence general-purpose models that can perform perception, control, and even non-robotic tasks [30, 31].
@@ 4.2.1 Robotics Action Generation Foundation Models
Language inputs have also been used for providing task specifications for end-to-end direct downstream control of robotics, forming language-conditioned imitation policy learning. Li et al. [250] use pre-trained language models to initialize policy networks that predict the next action. The control policy is fine-tuned via behavior cloning and active learning, to improve task completion. Directly using language inputs to generate robot actions for improving human-robot collaboration has also been proposed: the Language-Informed Latent Actions (LILA) framework [251] learns to use language to modulate the low-level controller, effectively using language to map a 7-DoF robotic arm output to a 2-DoF input. Further, Hu et al. [252] use multi-agent reinforcement learning to fine-tune an LLM policy to enable humans to specify what kind of strategies they expect from their AI partners. Building on the theme of instructing robots through language, Lynch et al. [253] use behavioral cloning on a dataset of hundreds of thousands of language-annotated trajectories to improve vision-language-motor skills in the real world. Though previous works mostly take a language-conditioned approach, RoboCat [195] trains imitation learning policy with just robotic data.
@@
As discussed in Section 2.1.3 (Preliminaries), offline RL has the potential to learn good policies without even interacting with the environment. With the availability of large-scale robotic datasets, offline RL starts to play an important role in developing effective RFMs. Early large-scale offline RL models such as QT-OPT [112] use a Q-learning-based approach in an offline manner to learn policy from robotics data which are collected by a robot farm. The successors of QT-OPT extend it to multitask learning by incorporating multi-task curriculum or predictive information [254, 255]. Recently, with the success of Transformer models, Q-learning based on transformer (Q-Transformer) also shows its potential [256]. PTR [257] is another promising work that adopts Conservative Q-Learning (CQL[258] in a multi-task learning setting. We look forward to seeing more RL-based robotic foundation models.
@@
Another direction of action foundation models involves vision or language pre-training  [259, 260, 261, 262, 263, 264, 28]. For example, inspired by the great generalization ability of self-supervised vision-based pre-training, MVP by Radosavovic et al. [261] trains visual representations on real-world images and videos from the internet and egocentric video datasets, via a masked autoencoder, and demonstrate the effectiveness of scaling visual pre-training for robot learning. Following this work, RPT [262] proposes Mask-pretraining with real robot trajectory data. VC-1 [264] actually did a comprehensive study on the effectiveness of vision-based pre-training on policy learning. We also recommend readers to learn more information about this problem from that paper. Despite the effectiveness of these visual pretraining methods, Hansen et al. [263] reexamined some of these methods and discovered significant domain gaps and propose a learning from scratch approach which remains competitive. This provides us a new perspective to think about visual pretraining in robotics.
@@ 4.2.2 General-purpose Robotics Foundation Models
4.3 How Foundation Models Can Help Solve Robotics Challenges
Modules Foundation Models Generalization 3.1 Data 3.2 Model 3.3 Task Specification 3.4 Uncertainty 3.5
Perception 2.1.1 VFM 2.2.1 Conceptgraphs [270] Conceptgraphs [270] - - -
VGM 2.2.2 -
@@ 4.3 How Foundation Models Can Help Solve Robotics Challenges
Task Planning and Action Generation  2.1.2 and 2.1.3 LLM 2.2.3 SayCan [24]
@@
Table 1: A summary of foundation models solving robotic challenges. Here we only list part of the works due to space limit. We find that uncertainly and safety are still largely unexplored.
@@
Foundation Models to Relieve the Requirement of Models As discussed in Section 3.3, building or learning a model—either a map of the environment, a world model, or an environmental dynamics model—is vital for solving robotic problems, especially in motion-planning and control. However, the powerful few/zero-shot generalization ability that foundation models present may break that requirement. This includes using LLMs to generate task plans [24], using RFMs to learn model-free end-to-end control policies [27, 256], etc.
Foundation Models for Task-Specification Task-specifications as language prompts [27, 28, 24], goal images [181, 272], videos of humans demonstrating the task [273, 274], rewards [26, 182], rough scratch of trajectory [239], policy sketches [275], and hand-drawn images [276] have allowed goal specifications in a more natural, human-like format. Multimodal foundation models allow users to not only specify the goal but also help resolve ambiguities via dialogue. Recent work in understanding trust and intent recognition within the human-robot interaction domain has opened up newer paradigms in our understanding of how humans use explicit and implicit cues to convey task-specifications. While significant progress has been made, recent work in prompt engineering for LLMs implies that even with a single modality, it is challenging to generate relevant outputs. Vision-Language Models are proven to be especially good at task-specification, showing potential for resolving this problem in robotics. Extending the idea of vision-language-based task-specifications, Cui et al. [181] explore methods to achieve multi-modal task specification using more natural inputs like images obtained from the internet. Brohan et al. [27] explores this idea of zero-shot transfer from task-agnostic data further, by providing a novel model class that exhibits promising scalable model properties. The model encodes high-dimensional inputs and outputs, including camera images, instructions, and motor commands into compact token representations to enable real-time control of mobile manipulators.
Foundation Models for Uncertainty and Safety Though being a critical problem in robotics, uncertainty and safety using foundation models for robotics is still underexplored. Existing works like KNOWNO [187] proposes a framework for measuring and aligning the uncertainty of LLM-based task planners. Recent advances in Chain-of-Thought prompting [277], open vocabulary learning [278], and hallucination recognition in LLMs [279] may open up newer avenues to address these challenges.
@@ 5.1 Datasets and Benchmarks
Relying solely on knowledge learned from language and vision datasets is limiting. Some concepts, like friction or weight, are not easily learned through these modalities alone, as suggested by Gao et al. [280] and Tatiya et al. [200] in their works on physically grounded VLMs. Therefore, in order to make robotic agents that can better understand the world, researchers are not just adapting foundational models from the language and vision domains; they are also advancing the development of large, diverse, and multimodal robotic datasets for training or fine-tuning these foundation models. This effort is now diverging into two directions: collecting data from the real world, versus collecting data from simulations and then transferring it to the real world. Each direction has its pros and cons. We will cover these datasets and simulations in the following paragraphs and discuss their respective advantages and disadvantages.
@@ 5.1.1 Real World Robotics Datasets
To overcome these limitations, some recent initiatives have made notable progress. For example, GNM [267] successfully integrated six different large-scale navigation datasets, utilizing a unified navigation interface based on waypoints. Furthermore, a recent collaborative effort among various laboratories called RT-X [176] has aimed to standardize data across different datasets, by using a 7-degree-of-freedom end-effector’s pose as a universal reference across different embodiments. This approach facilitates the joint use of diverse datasets and showed positive performance in cross-morphology transfer learning.
@@ 5.1.2 Robotics Simulators
Table 2: Tabletop Manipulation
Title Datasets & Simulation Real Robot
Number
of Tasks
Base Model SR SR Descriptions Frequency
RT-X
Kitchen
Manipulation [292],
Cable Routing [293],
NYU Door
Opening [294],
AUTOLab UR5 [295],
Robot Play [296],
Bridge [283],
RL Bench (RT-1) [27]
Google robot 160266 RT-2(PaLI)
30%
63%
92%
Bridge,
Small lab datasets,
Google robot
3-10Hz
RT-2
RL Bench (RT-1) [27]
Language-Table [253]
Google robot 480+ PaLI, PaLM-E
93%
62%
Seen tasks,
Unseen tasks
1-5Hz
RT-1[27] RL Bench (RT-1) [27] Google robot 744 RT-1
97%
76%
83%
59%
Seen Tasks,
Unseen Tasks,
With Distractors,
Novel background
3 Hz
GNFactor [198] RL Bench (RT-1) [27] XArm7 166
Stable Diffusion,
CLIP
32%
28%
Multi-task test,
Unseen tasks
Unspec.
MOO [249]
Self-created
robotic data
Google robot 106
Owl-ViT,
RT-1
92%
75%
Seen Tasks,
Unseen Tasks
Unspec.
PhysObjects [280] PhysObjects [280] Franka Panda 51
FlanT5-XXL,
GPT-4
88%
88%
PhysObjects Test Set,
Real scenes
open loop
Matcha [297] CoppeliaSim [298] NiCOL 50
ViLD,
text-davinci-003
91%
57%
Distinct descriptions,
Indistinct descriptions
open loop
Scalingup [236] Scalingup benchmark [236] UR5 arm 18 GPT-3 79% Mean Success 35 Hz
F3RM [197]
PyBullet [299] and
Self-created
robotic data
Franka Panda 18 CLIP
78%
62%
Grasp-Place,
Object Manip.
Unspec.
VIMA [272] VIMABench [272] None 17
pre-trained T5
tokenizer
81%
81%
79%
49%
L1,
L2,
L3,
L4
Unspec.
Instruct2Act [300] VIMABench [272] None 17
CLIP,
SAM,
text-davinci-003,
LLaMA
84% Mean Success open loop
VoxPoser [227]
Saipen [301]
Where2act [302]
Franka Emika 13
GPT-4,
SAM,
Owl-ViT
88%
70%
67%
Static environments,
With disturbances,
Unseen semarios
5HZ
Text2Motion [220] TableEnv [220] Franka Panda 6 text-davinci-003 82% Mean Success open loop
GenSim [222]
PyBullet [299] and
Self-created
robotic data
XArm7 100+ GPT-4
53.3%
68.8%
50 Tasks
70 Tasks
Unspec.
Table 3: Dexterous Manipulation
Title Datasets & Simulation Real Robot
Number
of Tasks
Base Model SR SR Descriptions Frequency
RoboCat [195]
RGB-Stacking
Benchmark [303]
Sawyer 7-DoF,
Panda 7-DoF,
KUKA 14-DoF
253 RoboCat
82%,
74%,
86%
Sawyer in sim,
Panda in sim,
KUKA in real
10 Hz
20 Hz
Table 4: Mobile Manipulation
Title Datasets & Simulation Real Robot
Number
of Tasks
Base Model SR SR Descriptions Frequency
LLaRP [304]
Language
Rearrangement [304]
None 1000 LLaMA-7B
42%
0%
8%
39%
Total,
Multiple objects,
Spatial Reasoning,
Conditional instuct.
open loop
Code-As-Policies [216]
Customized
RoboCodeGen [216],
HumanEval [305]
Eveyday Robots,
UR5 Robot arm
214
Codex,
GPT-3
95%
96%
RoboCodeGen,
HumanEval
open loop
InnerMonlogue [248] Ravens [306];
UR5e robot,
Google robot
130
ViLD
GPT-3.5
51%
50%
90%
60%
Seen Ravens tasks,
Unseen Ravens tasks,
Real robot arm,
Google robot
Unspec.
LIV [307]
MetaWorld [308]
FrankaKitchen [260]
Franka panda 114 LIV
30%
30%
45%
FrankaKitchen,
MetaWorld,
Real robots
15hz
SayCan [24]
proprietary simulator
and self-created
robotic data
Google robot 101
PaLM,
Flan
74% Mean Success open loop
PaLM-E [31] Lang-table [253]
Google robot,
xArm6
100 PaLM-E
83%
76%
52%
91%
Grasping,
Stacking,
Lang-table tasks,
Google robot
1-5Hz
TidyBot [309] TidyBot [309]
TidyBot
96
ViLD,
CLIP ,
text-davinci003
91%
83%
Sim,
Real
open loop
LLM-Grop [310] Gazebo [311]
Segway base
+ UR5e arm
8 GPT-3 4.08
Average human
rating(1-5)
open loop
LLM+P [247]
International Planning
Competition [223]
None 7 GPT-4
20%
90%
0%
95%
85%
20%
10%
Barman,
BLOCKSWORLD,
FLOORTILE,
GRIPPERS,
STORAGE,
TERMES,
TYREWORLD
open loop
HomeRobot [205] Habitat [312]
Hello-Robot Stretch
8 - 20%
RL baseline
closed loop
Table 5: Navigation
Title Datasets & Simulation Real Robot
Number
of Tasks
Base Model SR SR Descriptions
Frequency
LM-Nav [204]
Self-created datasets
Clearpath
Jackal UGV
20
GPT-3
80%
Mean Success
open loop
CLIP-Fields [202]
Habitat-Matterport
3D Semantic [312]
Stretch
Robot
14
CLIP
Detic
79%
Mean Success
open loop
GNM [267]
GNM datasets [267]
LoCoBot
Vizbot
DJI Tello
Jackal UGV
3
GNM
96%
Mean Success
Unspec.
Table 6: locomotion
Title Datasets & Simulation Real Robot
Number
of Tasks
Base Model SR SR Descriptions
Frequency
SayTap [26]
IsaacGym [171]
Unitree A1
30
GPT-4
97%
Mean Success
openloop
Prompt2Walk [229]
Mujoco [170]
None
1
GPT-4
80%
Mean Success
10 Hz
Table 7: Multi-Tasks
Title Datasets & Simulation Real Robot
Number
of Tasks
Base Model SR SR Descriptions Frequency
Gato [30]
Meta-World [308],
Sokoban [313],
BabyAI [314],
Procgen
Benchmark [315],
Arcade Learning
Environment [316],
DM Control
Suite [317]
Sawyer arm 604 GPT-4
97%
68%
80%
68%
super-human
64%
Meta-World,
Sokoban,
BabyAI,
Procgen,
ALE Atari,
DM Control Suite
20 Hz
Grounded
Decoding [318]
2D Maze [319],
Ravens [306]
None 124
GPT-3.5
PaLM
71%
46%
95%
51%
Ravens Seen Tasks,
Ravens Seen Tasks,
2D Maze,
Mobile Manipulation
open loop
Eureka [182]
Issac Gym [171],
Bidexterous
Manipulation [320]
None 29 GPT-4
55%
3.7
1
Bi-Dextrous,
Isaac Tasks –
(Human Score)
open loop
Lang2Reward [226] MuJoCo MPC [321] Google robot 17 GPT-4
95%
80%
Quadruped,
Dextrous manipulator
open loop
VC-1 [264] CortexBench [264] None 17 VC-1
59%
89%
67%
72%
60%
70%
63%
Adroit,
Meta-World,
DMControl,
Trifinger,
ObjectNav,
ImageNav,
Mobile Pick
open loop

### v3

@@ 4.1.3 LLMs and VLMs in Action Generation
4.1.4 Action Grounding
Traditionally, grounding refers to associating abstract and contextual meaning with sensory signals, such as identifying objects in images [135] or recognizing sounds in audio [136]. This form of ”direct grounding” in perceptual modalities has been extensively studied. However, in robotics, grounding extends to a crucial additional dimension: associating abstract concepts with executable actions. In other words, it involves translating the output of foundational models into meaningful changes in the environment. Thus, in this survey, we focus on action grounding.
Action grounding is just as vital as sensory grounding for robotics. For robots to interact with the physical world, foundational models require a bridge that maps their outputs to actionable behaviors, either explicitly or implicitly.
The initial approaches [26, 48, 130, 76] that tried to explicitly map foundation models’ outputs or representations to robot actions are fairly straightforward. The foundational models used are mainly large language models, and, like traditional robotics approaches, they adopt a modular concept that breaks down the robot’s preliminary skills. These skills are obtained either through pre-trained policies or pre-programmed scripts, and the entire skill library is described to the foundation models in natural language, allowing LLMs to output action plans using these skills. Recent progress in robotic grasping foundation models like AnyGrasp [137] and other open knowledge models like CLIP [23] and SAM [22] makes this approach more attractable. However, there are still two bottlenecks that it needs to solve: first, the limitations of natural language as a general interface, since the granularity that natural language can describe is limited. For example, if you were to use language to describe how to play Cat’s Cradle with a dexterous robot hand, language would likely struggle to accurately describe it. The second limitation is that, if we want the model to generalize to different tasks, more pre-trained or pre-programmed skills are needed, making it difficult to scale.
@@ 4.1.4 Action Grounding
Therefore, researchers are exploring different ways to address these issues. Code as an interface alleviates some limitations of natural language granularity [122, 121], as it can directly describe an object’s x, y, z position for grasping. Other novel interfaces try to involve directly anchoring foundational models to output joint torques, circumventing intermediary interfaces such as text tokens. A notable example is Gato [32], which directly maps the model’s output to the Atari game action space and robotic arm joint space. Gato dynamically decides its output format—be it text, joint torques, button presses, or other tokens—based on the context of the task at hand. Another related development is RT-2 [30], which, despite specifying the end-effector space in textual form, is capable of directly generating executable commands for robotic manipulator operation. SayTap [28] has experimented with directly generating binary foot patterns to control quadruped locomotion tasks.
Building on this idea of grounding models to real-world action spaces, approaches such as CLIP-Fields [94], VLMap [93], and NLMaps [27] project CLIP visual and semantic label representations directly onto 3D point clouds. By aligning semantic understanding with spatial information, these methods create more interpretable 3D maps for robotic applications. Going beyond explicit 3D mapping, GLAM [138] uses reinforcement learning to ground language models through interactions with the environment, demonstrating that LLMs can function effectively as RL agents in textual environments. Moreover, AutoRT [139] acts as an orchestrator for task performance by a fleet of robots, leveraging vision-language models [140, 141] for scene description to prompt LLMs for task generation.
Additionally, researchers are trying different interfaces to map the output of the foundation model to actions. Voxposer [129] maps the output to a 3D value map and then uses classical optimization techniques to calculate robot actions. This approach adds more robustness to disturbances in the pipeline and allows the model to manipulate trajectories to add safety constraints and avoid obstacles in real time. In addition to value maps, using constraints is also a popular approach to achieve grounding in more diverse tasks. Geometric constraints are used to compute feasible trajectories that avoid obstacles and achieve goals, while contact constraints are used to plan forceful or contact-rich behaviors. ReKep uses keypoint relationship constraints to represent both geometric and contact constraints as Python functions that map a set of keypoints to a numerical cost. Each keypoint is task-specific and represents a semantically meaningful 3D point in the scene.
While several grounding-to-action techniques are discussed above, there is no definitive answer as to which approach offers the most optimal solution. Pretrained skill libraries provide high levels of dexterity and precision in task execution, but at the expense of task diversity. Conversely, map- or constraint-based grounding techniques offer greater task flexibility but have primarily been demonstrated on simpler 2D gripper pick-and-place tasks, with no tests on more dexterous, complex actions. If we view grounding to action as a spectrum, the ideal interface should balance both task diversity and task complexity, excelling not only in the breadth of tasks it can perform but also in the intricacy of those tasks.
@@ 4.2 Robotics Foundation Models (RFMs)
With the increasing amount of robotics datasets, containing state-action pairs from real robots, the class of Robotics Foundation Models (RFMs) have likewise become increasingly viable [30, 67, 31, 69, 155, 156]. These models are characterized by the use of robotics data to train them, to solve robotics tasks. In this subsection, we summarize and discuss different types of RFMs. We will first introduce RFMs that can perform tasks according to a single robotic capability (e.g., perception, planning, and control), which is defined as single-purpose Robotic Foundation Models. For example, an RFM that can generate low-level actions to control the robot, or a model that can generate higher-level motion plans. We later introduce RFMs that can carry out tasks in multiple robotic modules, hence general-purpose models that can perform perception, control, and even non-robotic tasks [32, 33].
@@ 4.2.1 Robotics Action Generation Foundation Models
Imitation learning has been applied to robot control for a considerable period. Initially, methods primarily aimed at imitating a single skill. The concept of utilizing imitation learning to master multiple tasks, which is the focus of this review, emerged with the works of [158, 159]. This approach is known as one-shot imitation learning. Various conditions are employed to define the task, including robot goal images [158, 159, 160, 161, 162, 163, 164, 165], human goal images [16, 166, 167], language prompts [168, 169, 16], and task vectors [170], among others. A recent study also investigates the use of multi-modal input as a task descriptor [171]. These work represent early exploration in this direction. Lately, efforts have been made to scale this approach, emulating the success of large language models. Lynch et al.[172] employed behavioral cloning on a dataset of hundreds of thousands of language-annotated trajectories to learn tabletop block rearrangement in the real world. Brohan et al.[29] trained a large transformer model on over 130K tabletop rearrangement episodes with a more realistic setup and diverse objects, demonstrating the robustness of imitation learning algorithms on some unseen tasks. RoboCat[92] and RT-X[67] trained a single model with data from multiple embodiments, showing some generalization capability to unseen embodiments during testing. Overall, these efforts have primarily shown success in pick-and-place tasks or variants thereof, indicating significant potential for further exploration.
@@
With the availability of large-scale robotic datasets, offline RL starts to play an important role in developing effective RFMs. Early large-scale offline RL models such as QT-OPT [58] use a Q-learning-based approach in an offline manner to learn policy from robotics data which are collected by a robot farm. The successors of QT-OPT extend it to multitask learning by incorporating multi-task curriculum or predictive information [15, 173, 174]. Recently, with the success of Transformer models, Q-learning based on transformer (Q-Transformer) also shows its potential [175]. PTR [176] is another promising work that adopts Conservative Q-Learning (CQL[177] in a multi-task learning setting. When it comes to online RL, due to the large exploration space, most methods trained a visual policy in the simulation [178, 179] and then transferred the model to real world [180]. Still, this field is challenging, and we look forward to seeing more RL-based robotic foundation models.
@@
Another direction of action foundation models involves vision or language pre-training  [181, 182, 183, 184, 185, 186, 187, 30]. For example, inspired by the great generalization ability of self-supervised vision-based pre-training, MVP by Radosavovic et al. [184] trains visual representations on real-world images and videos from the internet and egocentric video datasets, via a masked autoencoder, and demonstrated the effectiveness of scaling visual pre-training for robot learning. Following this work, RPT [185] proposes mask-pretraining with real robot trajectory data. VC-1 [187] did a comprehensive study on the effectiveness of vision-based pre-training on policy learning. Despite the effectiveness of these visual pretraining methods, [186] reexamined some of these methods and discovered significant domain gaps, thus proposed learning-from-scratch approach. This provides us new perspective to think about visual pretraining in robotics.
@@
Using a similar approach as in pretraining with vision or language modality, we also see self-supervised pretraining with less-explored modalities such as audio  [189] and tactile sensing [190].
@@ 4.2.2 General-purpose Robotics Foundation Models
Modules Foundation Models Challenges
Generalization 3.1 Data 3.2 Model 3.3 Task Specification 3.4 Uncertainty 3.5
Perception VFM Conceptgraphs [194] Conceptgraphs [194] - - -
VGM -
@@
Task Planning and Action Generation LLM SayCan [26]
@@
Table 1: A quick summary of foundation models solving robotic challenges. Here we only list part of the works due to space limit. We find that uncertainly and safety are still largely unexplored.
4.3 How do Foundation Models Help Solve Robotics Challenges
@@
Foundation Models to Relieve the Requirement of Models
As discussed in Section 3.3, building or learning a model—either a map of the environment, a world model, or an environmental dynamics model—is vital for solving robotic problems, especially in motion-planning and control. However, the powerful few/zero-shot generalization ability that foundation models present may break that requirement. This includes using LLMs to generate task plans [26], using RFMs to learn model-free end-to-end control policies [29, 175], etc.
Foundation Models for Task-Specification
Task-specifications as language prompts [29, 30, 26], goal images [77, 171], videos of humans demonstrating the task [199, 200], rewards [28, 78], rough scratch of trajectory [145], policy sketches [201], and hand-drawn images [202] have allowed goal specifications in a more natural, human-like format. Multimodal foundation models allow users to not only specify the goal but also help resolve ambiguities via dialogue. Recent work in understanding trust and intent recognition within the human-robot interaction domain has opened up newer paradigms in our understanding of how humans use explicit and implicit cues to convey task-specifications. While significant progress has been made, recent work in prompt engineering for LLMs implies that even with a single modality, it is challenging to generate relevant outputs. Vision-Language Models are proven to be especially good at task-specification, showing potential for resolving this problem in robotics. Extending the idea of vision-language-based task-specifications, explore methods to achieve multi-modal task specification using more natural inputs like images obtained from the internet  [29] explores this idea of zero-shot transfer from task-agnostic data further, by providing a novel model class that exhibits promising scalable model properties. The model encodes high-dimensional inputs and outputs, including camera images, instructions, and motor commands into compact token representations to enable real-time control of mobile manipulators.
Foundation Models for Uncertainty and Safety
Though being a critical problem in robotics, uncertainty and safety using foundation models for robotics is still underexplored. Existing works like KNOWNO [83] proposes a framework for measuring and aligning the uncertainty of LLM-based task planners. Recent advances in Chain-of-Thought prompting [203], open vocabulary learning [204], and hallucination recognition in LLMs [205] may open up newer avenues to address these challenges.
@@ 5.1 Datasets and Benchmarks
Relying solely on knowledge learned from language and vision datasets is limiting. Some concepts, like friction or weight, are not easily learned through these modalities alone, as suggested by [206] and [115] in their works on physically grounded VLMs. Therefore, in order to make robotic agents that can better understand the world, researchers are not just adapting foundational models from the language and vision domains; they are also advancing the development of large, diverse, and multimodal robotic datasets for training or fine-tuning these foundation models. This effort is now diverging into two directions: collecting data from the real world, versus collecting data from simulations and then transferring it to the real world. Each direction has its pros and cons. We will cover these datasets and simulations in the following paragraphs and discuss their respective advantages and disadvantages.
@@ 5.1.1 Real World Robotics Datasets
To overcome these limitations, some recent initiatives have made notable progress. For example, GNM [191] successfully integrated six different large-scale navigation datasets, utilizing a unified navigation interface based on waypoints. A recent collaborative effort among various laboratories called RT-X [67] has aimed to standardize data across different datasets, by using a 7-degree-of-freedom end-effector’s pose as a universal reference across different embodiments. To offer a more comprehensive data source, RH20T [68] collected over 110,000 manipulation episodes across 7 embodiments, covering more than 140 diverse, contact-rich skills. The modalities include RGB, depth, end-link force-torque, tactile, proprioception, audio and language instruction. All sensors are well-calibrated and synchronized.